the sum of two numbers is 4. find the minimum value of the sum of their cubes.


1
Expert's answer
2022-03-19T02:35:38-0400


Solution


Let the two numbers are "x" and "y"


Then, according to the given condition


"x+y=4" ... (1)


And the sum of their cubes is


"S=x^3+y^3" ... (2)


From (1) we can write "y=4-x", replace in (2)


"S=x^3+(4-x)^3"


We need to find the minimum sum, therefore, we minimize it using the first derivative


"\frac{dS}{dx}=3x^2+3(4-x)^2(-1)"


Replacing this first derivative equal to zero,


"\frac{dS}{dx}=0"


"3x^2+3(4-x)^2(-1)=0"


"3x^2-3(16-8x+x^2)=0"


"3x^2-48+24x-3x^2=0"


"-48+24x=0"


"x=\frac{48}{24}=2"


Which is the first number for which the sum is minimum


And therefore, the second number is


"y=4-x"


"y=4-2=2"


Hence the two numbers are "2" and "2"